Superior Company, a manufacturer of cedar picnic tables, had the following data for 2017: INSTRUCTIONS: Using the information provided answer the following questions. Show your work. Sales 2,400 units $40 per unit $15 per unit $19,500 Sales price Variable costs Fixed costs 1) Prepare a CVP Income Statement for 2017 (5 pt) 2) What is the contribution margin ratio (1 pt)? 3) What is the break-even point in dollars (1 pt)? 4) What is the margin of safety in dollars (1 pt)? 5) If the company wishes to increase its total dollar contribution margin by 40% in 2018, by how much will it need to increase sales if all other factors remain constant (2 pt)?
Answer #1

1)    Superior Company

CVP Income Statement

For the year ended 2017

Particulars Total Per Unit
Sales (2400 * 40) 96000 40
Variable Costs (2400 * 15) 36000 15  
Contribution Margin 60000 25
Fixed Costs 19500
Net Income 40500

2)  Contribution Margin Ration = Contribution margin / Sales *100

= 60000 / 96000 * 100

= 62.5 %

3) Break-even in Dollars = Fixed Costs / Contribution margin ration

= 19500 / 62.5% = 31200

4) Margin of safety in Dollars = Sales - Break-even sales

= 96000 - 31200 = 64800

5)      Contribution margin * (percentage of increase) 40%

= 60000 * 1.4 = 84000

increased contribution margin / contribution margin ratio

= 84000 / 62.5% = 134400 (Sales)